OpenStudy (anonymous):

how can you find the area of a trapezoid using the formula 1/2(B+b)h. (B) Base equals 5' and the (b) base equals 10'. the (h) height from base to base equals 12'. please show work.thanks!

OpenStudy (across):

All you have to do is plug in the given variables.

OpenStudy (across):

\[\frac{1}{2}(B+b)h\]\[\frac{1}{2}(5+10)12\]\[\frac{1}{2}(15)12\]\[\frac{1}{2}180\]\[90\]